# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: Present address: Nord Quantique, Sherbrooke, QC J1J 2E2, Canada

### v1

@@ V.2.4 Adjusted Kerr
Upon close inspection of Fig. 3, a slight deviation between the theoretical location of the resonances and the experimental observations comes to light. This deviation seems to increase with $\epsilon_{1}/K$, suggesting there could be a constant scaling factor between the independently calibrated $\epsilon_{1}/K$ and the actual value. Indeed, allowing for a small rescaling by $5\%$ gives great agreement between theory and experiment, as demonstrated in Fig. 15. While this is not within the calibration uncertainties (Fig. 7), past experiments have shown that the driven Kerr under the squeezing drive can be slightly renormalized compared to the spectroscopic measurement of Kerr [15, 24]. In Fig. 16 B, we show the Kerr adjusted experimental data from Fig. 4 B.
Figure 17: Numerical quantum treatment and analysis. A) Heatmap of $\mathcal{J}$ (numerical quantum treatment) with the semiclassical prediction on top (see Fig. 4 C). B) Plot of $\mathcal{J}$ along the red dashed line in A parameterised by $\epsilon_{2}/K$. Vertical dashed lines indicate the location of the singularities of $\mathcal{J}$ which happen very close to the triple blue-orange intersections. C) Plot of $\mathcal{J}$ along the horizontal black dashed line at $\epsilon_{2}/K=10$ in A parameterised by $\epsilon_{1}/K$. Vertical dashed lines indicate the resonance condition from Eq. 4. The different semiclassical predictions agree well with the full quantum Hamiltonian calculation.
Figure 18: Experimental lifetimes of Fig. 4 A along each resonance parabola as parameterised by $\epsilon_{1}/K$. A-D) show the the four parabolas for $n=1,2,3,4$ as defined by resonance condition Eq. 4. The orange curves are guides to the eye (a box average).
@@ V.2.5 Lifetime along resonances
In Fig. 18 we display the lifetime along the first four resonances/parabolas in Fig. 4 B. As per resonance condition Eq. 4, an increase in $\epsilon_{1}/K$ corresponds to an increase in $\epsilon_{2}/K$ when following the parabolas. Therefore, we expect the lifetime to increase along the parabolas, which is what we observe in Fig. 18. The step like behaviour is a consequence of activation assisted tunneling close to the separatrix between the wells (see [15]).
Full quantum Hamiltonian treatment
@@
$\mathcal{J}\doteq\sum_{n=0}^{N_{\mathrm{cff}}}\frac{1}{\log\left(E_{n+1}-E_{n}\right)},$ (6)
@@
By construction, $\mathcal{J}$ diverges when a new pair of levels in $\{E_{0},E_{1},\cdots,E_{N_{\mathrm{cff}}}\}$ are closing. To see this, consider the scenario where the gap between the pair $E_{0}$ and $E_{1}$ and the gap between the pair $E_{2}$ and $E_{3}$ are closed, and the rest of the levels are not anti-crossing. In this situation, $\mathcal{J}$ is roughly constant since $\frac{1}{\log\left(E_{n+1}-E_{n}\right)}$ is zero for the two closed pairs and finite for $E_{2}-E_{1}$ and the rest of the levels. Consider now that $E_{4}$ and $E_{5}$ are interacting, starting to close their gap as in Fig. 2 D. These interacting levels will close fast and at some point have an energy difference $E_{4}-E_{5}=1$ where $\frac{1}{\log\left(E_{n+1}-E_{n}\right)}$ will diverge. Also, this approach to $E_{4}-E_{5}=1$ happens differently depending on how one approaches the resonance in parameter space, approaching $1$ from the left or from the right in $\epsilon_{2}$ makes $\mathcal{J}$ abruptly change sign near a new anti-crossing. Observe the close agreement between the semiclassical treatment, the resonance condition Eq. 4, and $\mathcal{J}$.
V.3 Effects beyond the RWA model
@@
Figure 19: Unexplained resonance-like features. A) Coherent state lifetime as a function of photon number $\epsilon_{2}/K$ (at $\epsilon_{1}=0$). The blue and red curve are the same measurement repeated after a 24h time difference. The frequency was recalibrated to account for a small drift. The lifetime saturates at around 300 µs. For large photon numbers $\epsilon_{2}/K$, resonance-like structures appear. The location of these is stable within the 24h time difference. B) The parametric oscillator is initialized in one of the wells and after 24 µs the remaining population in this well is measured. This is a proxy for the coherent state lifetime and thus allows us to identify resonances like in A, which emerge as a low remaining population. This is measured as a function of photon number $\epsilon_{2}/K$ and detuning $\Delta=\frac{\omega_{2}}{2}-\omega_{a}$ between the first subharmonic of the squeezing drive and the SNAIL transmon resonance frequency. At $\epsilon_{2}/K=12$ and $\Delta/K=1$ a crossing between resonances is visible. Another feature occurs at $\epsilon_{2}/K=17$ and $\Delta/K=2$, where the line of a resonance splits into two.
@@
Figure 20: Emergence of resonance-like structures depending on different parameters. A-C) Relative phase $\phi=0\degree$ between linear drive and squeezing drive. A) Equienergy contours of the classical RWA parametric oscillator Hamiltonian with linear drive. The linear drive raises the left well and breaks the symmetry between the two wells within the RWA. B) Same experiment as in Fig. 3 A, now at $\epsilon_{2}/K=18$. A “forest” of resonances is now visible besides the RWA resonances from condition Eq. 4. C) Changing the photon number to $\epsilon_{2}/K=16$ changes the resonance pattern. Note the different widths. D-E) Relative phase between linear and squeezing drive of $\phi=90\degree$. D) Equienergy surfaces in the presence of the linear drive within the RWA, showing that the symmetry between wells is preserved and a small deformation in the $p$ direction is visible. E) Same measurement as in B but with $90\degree$ shifted phase $\phi$. This change in phase is sufficient to remove all of the resonances, while the amplitude of the linear drive remains high. F) Extension of data in Fig. 4 B. Now the photon number extends beyond $\epsilon_{2}/K=12$ up to $\epsilon_{2}/K=18$. The data shown in B, C show up as flecks of short lifetime (blue) that are not captured by the RWA.
@@
In this section, we present novel experimental data taken for $\epsilon_{2}/K>12$ where the static effective description used fails, qualitatively, to describe experiment. In Fig. 19 A, we present the coherent state lifetime up to $\epsilon_{2}/K\approx 17$, for $\epsilon_{1}=0$. The lifetime increases in a step-like fashion, as predicted by the static effective model [29, 28] until it starts decreasing. This decrease may be captured by a more elaborated static effective treatment [30]. However, resonant-like drops of the lifetime appear too, which challenge the static effective treatment, even if they can also be captured in principle [44]. Time-dependent Floquet simulation of the driven system (not shown, see [21]) suggest that these resonances may be contained within a single-mode treatment of the nonlinear transmon oscillator. Note that, since these experiments were conducted with $\epsilon_{1}=0$, these resonances are of a different yet similar nature from the ones presented in the main text. Below, we present and discuss control experiments supporting the viewpoint that these effects belong to the nonlinear physics of single-mode driven systems.
@@
As a control experiment to learn more about these resonances, we measure their temporal stability by repeating the same measurement again after 24h (see orange and blue graphs in Fig. 19 A). This provides important information since the parameters of spurious two-level-systems (TLS) [48, 49, 50] coupled to transmons are known to fluctuate in time, and TLS are a plausible candidate for this resonant-like lifetime drops at strong drives. We re-calibrated the qubit frequency and observed good agreement between the two measurements, with the resonances being located at the same positions. This data suggests that the resonances are stable in time and are therefore likely associated to stable electromagnetic transitions and not fluctuating TLSs.
@@
To learn more, we study these resonances as a function of the detuning $\Delta=\frac{\omega_{2}}{2}-\omega_{a}$ between the squeezing drive and the resonance frequency of the SNAIL transmon. The system is initialized in one well and after 24 µs the remaining population is measured. This is a proxy for the coherent states lifetime and thus resonances show up through low remaining population. Fig. 19 B depicts this lifetime proxy as function of detuning and photon number. This data contains resonances moving in different directions. Other features can be seen, for example, around $\epsilon_{2}/K=12$ and $\Delta/K=1$: here two resonances cross (i.e., no avoided crossing), which could provides a bound to the possible coupling between the involved modes. Lastly, we point out a feature around $\epsilon_{2}/K=17$ and $\Delta/K=2$, where one of the resonances seems to split up into two.
@@
Moving beyond the case of $\epsilon_{1}=0$, we now study the impact of an additional drive on the resonances. In Fig. 20 A, we show the equienergy contours of the effective Hamiltonian for $\epsilon_{1}\neq 0$ and $\phi=0$, which creates an asymmetry between the two wells (see also Fig. 2 C). This picture is valid under the RWA, and we have observed its conspicuous failure for $\epsilon_{2}/K\gg 1$. In Fig. 20 B, for $\epsilon_{2}/K=18$ we observe a “forest” of resonances, very different from the case of smaller $\epsilon_{2}/K$ (see Fig. 3). These resonances are not explained by resonant tunneling in the RWA potential and described by Eq. 4, but are instead reminiscent to the onset of chaotic behavior [51, 21]. Changing the photon number (see Fig. 20 C) leads to a change in the forest of resonances. Similar resonances have been observed during readout of transmon qubits. Here a possible explanation is that the AC Stark shift, induced by the strong readout drive, tunes the qubit into resonance with lossy modes like TLS [47, 52, 46]. To test this hypothesis, we change the phase $\phi$ of our strong linear drive. For $\phi=90\degree$, the RWA double-well stays symmetric (Fig. 20 D). In Fig. 20 E, we show the same measurement as in Fig. 20 B, (same value of $\epsilon_{1}$), but with a different relative phase between the drives. The absence of resonances in Fig. 20 E suggests that the effect is beyond a Stark shift into lossy modes. This phase dependence of the spurious resonances has not been observed before. Finally, Fig. 20 F shows an extended data set (compare to Fig. 4), where the breakdown of the RWA description is self-evident for $\epsilon_{2}/K>12$.
@@
Our data set, presenting resonances at different locations in parameter space and of different widths, as well as their dependence on the parametric drive frequency, the parametric drive amplitude, the linear drive amplitude, the relative phase between the linear drive and the parametric drive, and their stability over a period of 24 hs, will guide research and lead to better theoretical tools to understand and design parametric processes. We also expect that this data will unlock new tools and proposals to study nonlinear driven quantum systems and quantum chaos in unexplored regimes [55, 56, 57].

### v2

@@ V.2.4 Adjusted Kerr
Upon close inspection of Fig. 3, a slight deviation between the theoretical location of the resonances and the experimental observations comes to light. The theory in Fig. 3 has no free parameters.
This deviation seems to increase with $\epsilon_{1}/K$, suggesting there could be a constant scaling factor between the independently calibrated $\epsilon_{1}/K$ and the actual value. Indeed, allowing for a small rescaling by $5\%$ gives great agreement between theory and experiment, as demonstrated in Fig. 15. While this is not within the calibration uncertainties (Fig. 7), past experiments have shown that the driven Kerr under the squeezing drive can be slightly renormalized compared to the spectroscopic measurement of Kerr [16, 25]. In Fig. 16 B, we show the Kerr adjusted experimental data from Fig. 4 B.
Figure 17: Numerical quantum treatment and analysis. A) Heatmap of $\mathcal{J}$ (numerical quantum treatment) with the semiclassical prediction on top (see Fig. 4 C). B) Plot of $\mathcal{J}$ along the green dashed line in A parameterised by $\epsilon_{2}/K$. Vertical black dashed lines indicate the location of the singularities of $\mathcal{J}$ which happen very close to the triple blue-orange intersections. C) Plot of $\mathcal{J}$ along the horizontal blue dashed line at $\epsilon_{2}/K=10$ in A parameterised by $\epsilon_{1}/K$. Vertical black dashed lines indicate the resonance condition from Eq. 4. The different semiclassical predictions agree well with the full quantum Hamiltonian calculation.
Figure 18: Experimental lifetimes of Fig. 4 A along each resonance parabola as parameterised by $\epsilon_{1}/K$. A-D) show the the four parabolas for $n=1,2,3,4$ as defined by resonance condition Eq. 4. The orange curves are guides to the eye (a moving average).
@@ V.2.5 Lifetime along resonances
In Fig. 18 we display the lifetime along the first four dashed parabolas of Fig. 4 B. As per resonance condition Eq. 4, an increase in $\epsilon_{1}/K$ corresponds to an increase in $\epsilon_{2}/K$ when following the parabolas. Therefore, we expect the lifetime to increase along the parabolas, which we observe in Fig. 18. The step like behaviour is a consequence of activation assisted tunneling close to the separatrix between the wells (see [16]).
V.3 Full quantum Hamiltonian treatment
@@
$\mathcal{J}\doteq\sum_{n=0}^{N_{\mathrm{cff}}}\Bigg{|}\frac{1}{\log\left(E_{n+1}-E_{n}\right)}\Bigg{|},$ (6)
@@
By construction, $\mathcal{J}$ diverges when a new pair of levels in $\{E_{0},E_{1},\cdots,E_{N_{\mathrm{cff}}}\}$ are closing. To see this, consider the scenario where the gap between the pair $E_{0}$ and $E_{1}$ and the gap between the pair $E_{2}$ and $E_{3}$ are closed, and the rest of the levels are not anti-crossing. In this situation, $\mathcal{J}$ is finite, since $\big{|}\frac{1}{\log\left(E_{n+1}-E_{n}\right)}\big{|}$ is zero for the two closed pairs and finite for $E_{2}-E_{1}$ and the rest of the levels (which are distant from one another). Consider now levels $E_{4}$ and $E_{5}$ interacting, starting to close their gap as in Fig. 2 D. These interacting levels will close and at some point have an energy difference $E_{4}-E_{5}=1$ where $\big{|}\frac{1}{\log\left(E_{n+1}-E_{n}\right)}\big{|}$ will diverge. In this way, $\mathcal{J}$ picks up poles near new anti-crossings reproducing Fig. 4 as it can be seen in Fig. 17 B and C in greater detail. Observe the close agreement between the semiclassical treatment, the resonance condition Eq. 4, and $\mathcal{J}$.
V.4 Effects beyond the RWA model
@@
Figure 19: Resonance-like features not explained by RWA model. A) Coherent state lifetime as a function of photon number $\epsilon_{2}/K$ (at $\epsilon_{1}=0$). The blue and red curves are the same measurement repeated after a 24 hours time difference. The frequency was recalibrated to account for a small drift. The lifetime saturates at around 300 µs. For large photon numbers $\epsilon_{2}/K$, resonance-like features appear. The location of these is stable within the 24 hours time difference. B) The parametric oscillator is initialized in one of the wells and after 24 µs the remaining population in this well is measured. This is a proxy for the coherent state lifetime and thus allows us to identify resonances like in A, which emerge as a low remaining population. This is measured as a function of photon number $\epsilon_{2}/K$ and detuning $\Delta=\frac{\omega_{2}}{2}-\omega_{a}$ between the first subharmonic of the squeezing drive and the SNAIL transmon resonance frequency. At $\epsilon_{2}/K=12$ and $\Delta/K=1$ a crossing between resonances is visible. Another feature occurs at $\epsilon_{2}/K=17$ and $\Delta/K=2$, where the line of a resonance splits into two.
@@
Figure 20: Emergence of resonance-like features depending on different parameters. A-C) Relative phase $\phi=0\degree$ between linear drive and squeezing drive. A) Equienergy contours of the classical RWA parametric oscillator Hamiltonian with linear drive. The linear drive raises the left well and breaks the symmetry between the two wells within the RWA. B) Same experiment as in Fig. 3 A, now at $\epsilon_{2}/K=18$. A “forest” of resonances is now visible besides the RWA resonances from condition Eq. 4. C) Changing the photon number to $\epsilon_{2}/K=16$ changes the resonance pattern. Note the different widths. D-E) Relative phase between linear and squeezing drive of $\phi=90\degree$. D) Equienergy surfaces in the presence of the linear drive within the RWA, showing that the symmetry between wells is preserved and a small deformation in the $p$ direction is visible. E) Same measurement as in B but with $90\degree$ shifted phase $\phi$. This change in phase is sufficient to remove all of the resonances, while the amplitude of both drives remains high. F) Extension of data in Fig. 4 B. Now the squeezing amplitude extends beyond $\epsilon_{2}/K=12$ up to $\epsilon_{2}/K=18$. The data in B, C shows up as flecks of short lifetime (blue) that are not captured by the RWA.
@@
In this section, we present experimental data taken for $\epsilon_{2}/K>12$ where the static effective (RWA) description used fails, qualitatively, to describe experiment. In Fig. 19 A, we present the coherent state lifetime up to $\epsilon_{2}/K\approx 17$, for $\epsilon_{1}=0$. The lifetime increases in a step-like fashion, as predicted by the static effective model [30, 29] until it starts decreasing. This decrease may be captured by a more elaborated static effective treatment [31]. However, resonant-like drops of the lifetime appear too, which challenge the static effective treatment, even if they can also be captured in principle [45]. Time-dependent Floquet simulation of the driven system (not shown, see [22]) suggest that these resonances may be explained by a single-mode treatment of the nonlinear transmon oscillator. Note that, since these experiments were conducted with $\epsilon_{1}=0$, these resonances are of a similar nature to the ones presented in the main text. Below, we present and discuss control experiments supporting the viewpoint that these effects belong to the nonlinear physics of single-mode driven systems.
@@
As a control experiment to learn more about these resonances, we measure their temporal stability by repeating the same measurement again after 24 hours (see orange and blue graphs in Fig. 19 A). This provides important information since the parameters of spurious two-level-systems (TLS) [49, 50, 51] coupled to transmons are known to fluctuate in this timescale, and TLS are a plausible candidate for this resonant-like lifetime drops at strong drives. We re-calibrated the qubit frequency and observed good agreement between the two measurements, with the resonances being located at the same positions. This data suggests that the resonances are stable in time and are therefore likely associated to stable electromagnetic transitions and not fluctuating TLSs.
@@
Next, we study these resonances as a function of the detuning $\Delta=\frac{\omega_{2}}{2}-\omega_{a}$ between the squeezing drive and the resonance frequency of the SNAIL transmon. The system is initialized in one well and after 24 µs the remaining population is measured. This is a proxy for the coherent states’ lifetime and thus resonances manifest themselves as a low remaining population. Fig. 19 B depicts this lifetime proxy as function of detuning and photon number. This data contains resonances moving in different directions. Other features can be seen, for example, around $\epsilon_{2}/K=12$ and $\Delta/K=1$: here two resonances cross (i.e., no avoided crossing), which could provide a bound to the possible coupling between the involved modes. Lastly, we point out a feature around $\epsilon_{2}/K=17$ and $\Delta/K=2$, where one of the resonances seems to split up into two.
@@
Moving beyond the case of $\epsilon_{1}=0$, we now study the impact of an additional drive on the resonances. In Fig. 20 A, we show the equienergy contours of the effective Hamiltonian for $\epsilon_{1}\neq 0$ and $\phi=0$, which creates an asymmetry between the two wells (see also Fig. 2 C). This picture is valid under the RWA, and we have observed its conspicuous failure for $\epsilon_{2}/K\gg 1$. In Fig. 20 B, for $\epsilon_{2}/K=18$ we observe a “forest” of resonances, very different from the case of smaller $\epsilon_{2}/K$ (see Fig. 3). These resonances are not explained by resonant tunneling in the RWA potential and described by Eq. 4, but are instead reminiscent to the onset of chaotic behavior [52, 22, 53]. Changing the photon number (see Fig. 20 C) leads to a change in the forest of resonances. Similar resonances have been observed during the readout of transmon qubits. Here a possible explanation is that the AC Stark shift, induced by the strong readout drive, tunes the qubit into resonance with lossy modes like TLS [48, 54, 47]. To test this hypothesis, we change the phase $\phi$ of our strong linear drive. For $\phi=90\degree$, the RWA double-well stays symmetric (Fig. 20 D). In Fig. 20 E, we show the same measurement as in Fig. 20 B, (same value of $\epsilon_{1}$), but with a different relative phase between the drives. The absence of resonances in Fig. 20 E suggests that the effect is beyond a Stark shift into lossy modes. This phase dependence of the spurious resonances has not been observed before. Finally, Fig. 20 F shows an extended data set (compare to Fig. 4), where the breakdown of the RWA description is self-evident for $\epsilon_{2}/K>12$.
@@
Our data set, presenting resonances at different locations in parameter space and of different widths, as well as their dependence on the parametric drive frequency, the parametric drive amplitude, the linear drive amplitude, the relative phase between the linear drive and the parametric drive, and their stability over a period of 24 hs, will guide research and lead to better theoretical tools to understand and design parametric processes. We also expect that this data will unlock new tools and proposals to study nonlinear driven quantum systems and quantum chaos in unexplored regimes [57, 58, 59, 53].